Solve the equation: 2/x-1 = 6x/x+4 Please show me how to solve this Thanks!
Answer by Joker(6)

Cross multiply the equation first so you get 2(x+4)=6x(x-1). Multiply everything out which gets you 2x+8=6x^2-6x. move everything to one side so it will equal to zero and then solve for x which looks like: 6x^2-8x-8=0 from here, you factor it into (6x+4)(x-2)=0 now set both parenthesis = to 0 so you get x = -2/3 and 2.